40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4663     Entry time: Mon May 9 09:37:51 2011
Author: valera 
Type: Update 
Category: PSL 
Subject: PSL and MC trends 

The attached plot shows 7 day trends of the MC and PMC power levels, PSL QPDs, and temperature. The MC stayed locked for ~40 hours over the weekend. The temperature swings were somewhat smaller over the past couple of days but one should remember to turn the PSL HEPA down after working on the table. Steve turned the HEPA flow from 100% down to 20% on Thursday and posted the reminder signs on the PSL enclosure.

Attachment 1: pslmcdrift2.pdf  197 kB  | Hide | Hide all
pslmcdrift2.pdf
ELOG V3.1.3-